OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

-2x - 6 = -2x - 2 - 4 combine like terms

OpenStudy (compassionate):

-2x -3 = -2x - 2 - 4 Transpose: -2x + 2x = 3 - 2 - 4 0 = 3 - 6 <-- Do the last part What happens when the x's cancel and become 0's?

hero (hero):

sorry divide everything by -2 to get x + 3 = x + 1 + 2

hero (hero):

Then you'll realize that x + 3 = x + 3

hero (hero):

But you won't be able to solve for x

hero (hero):

However, you can subtract x from both sides and subtract 3 from both sides to get 0 = 0 Then you'll know that the solutions are infinite.

hero (hero):

actually, @Compassionate made a mistake with his solution while trying to distribute the -2 on the left side.

OpenStudy (anonymous):

May you help me with this one? x = 0.8x + 72

hero (hero):

subtract 0.8x from both sides to end up with 0.2x = 72 Divide both sides by 0.2 to get x = 72/0.2 x = 360